Find both the x intercept and the y intercept of the line given by the equation
x+8y+5=0

Answer by nerdybill(6962)

To find the y-intercepts, set x=0 and solve for y:
x+8y+5=0
0+8y+5=0
8y=-5
y= -5/8
Therefore, the y-intercept is at (0, -5/8)
.
To find the x-intercepts, set y=0 and solve for x:
x+8y+5=0
x+8(0)+5=0
x+5=0
x = -5
Therefore, the x-intercept is at (-5, 0)